ROOM-TEMPERATURE CHIRPED-PULSE FOURIER TRANSFORM MICROWAVE (CP-FTMW) SPEC-TRUM OF 2-METHYLFURAN

\underline{IAN A. FINNERAN} and STEVEN T. SHIPMAN, Division of Natural Sciences, New College of Florida, Sarasota, FL 34243.

The microwave spectrum of 2-methylfuran has been measured at temperatures between 0 ^{∘}C and 50 ^{∘}C from 8.7–18.3 GHz with waveguide chirped-pulse Fourier transform microwave spectroscopy (CP-FTMW). Using the enhanced sensitivity of this technique relative to that of prior measurements from Stark-modulated instruments, we have been able to extend the assignments of the lowest energy A- and E-states from Norris and Krisher^{a} to include transitions up to J = 60 and Ka = 35. We will also report on our progress towards assigning higher-lying states and compare the fit results to those from ab initio calculations.

aNorris, W. and Krisher, L., J. Chem. Phys. 51, 403 (1969).